OpenStudy (anonymous):

You are pulling a sled and a rider with combined mass of 82 kg up a 6.5 degree slope at a steady speed. If the coefficient of kinetic friction between the sled and snow is 0.1, what is the tension in the rope?

OpenStudy (anonymous):



OpenStudy (anonymous):

So \[F-mg\sin{6.5}-f=ma\\F-mg\sin6.5-\mu N=ma \] Now,\[N-mg\cos6.5=0\\N=mg\cos6.5\] Substitute to prev eqtn:\[F-mg\sin6.5-\mu mg\cos6.5 =ma\\a=\frac{F-mg\sin6.5-\mu mg\cos6.5}{m} \]

OpenStudy (anonymous):

Now, from that equation, we see that we need to know a to find F (Tension of the rope). It means tension will vary depend on acceleration of the sled+rider

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what do I do next?

OpenStudy (anonymous):

The problem says the sled is pulled at a steady speed, so the acceleration = 0